Write an equation in point-slope form and slope intercept form:
Slope= -1, passing thru (-1/2, -2)

Answer by Cromlix(4381)

y - b = m(x - a)
y + 2 = -1(x + 1/2)
y = -x - 1/2 - 2
y = -x - 5/2
or
2y = -2x -5